What is the magnitude of the electric force of attraction between an iron nucleus

(q = −26e) and its innermost electron if the distance between them is 1.5 × 10−12m?


1
Expert's answer
2022-02-28T09:59:48-0500

Answer

the magnitude of the electric force

"F=\frac{KQq}{r^2}\\=
\frac{9*10^9*26*1.6*10^{-19}*1.6*10^{-19}}{1.5*1.5*10^{-24}}\\=0.00234N"